Cell Potential at Equilibrium For a single galvanic cell based on the (unbalanced) reaction: Ag+ (aq) + Zn(s) = Zn2+(aq) + AgPlease show all work step by step and final answer.
Answer #1

At equilibrium every cell has 0 potential value

But I will show you E° value for your cell

2 Ag+(aq) + 2e = 2 Ag(s). E° = 0.8

Zn(s) = Zn2+(aq) + 2e E°= 0.76

E° = 0.76 + 0.8

= 1.56 V

If you have any questions please comment

If you satisfied with the solution please rate it thanks